### v2

# NAAQA: A Neural Architecture for Acoustic Question Answering

Abstract
The goal of the Acoustic Question Answering (AQA) task is to answer a free-form text question about the content of an acoustic scene. It was inspired by the Visual Question Answering (VQA) task. In this paper, based on the previously introduced CLEAR dataset, we propose a new benchmark for AQA, namely CLEAR2, that emphasizes the specific challenges of acoustic inputs. These include handling of variable duration scenes, and scenes built with elementary sounds that differ between training and test set. We also introduce NAAQA, a neural architecture that leverages specific properties of acoustic inputs. The use of 1D convolutions in time and frequency to process 2D spectro-temporal representations of acoustic content shows promising results and enables reductions in model complexity. We show that time coordinate maps augment temporal localization capabilities which enhance performance of the network by $\sim$17 percentage points. On the other hand, frequency coordinate maps have little influence on this task. NAAQA achieves 79.5% of accuracy on the AQA task with $\sim$4 times fewer parameters than the previously explored VQA model. We evaluate the perfomance of NAAQA on an independent data set reconstructed from DAQA. We also test the addition of a MALiMo module in our model on both CLEAR2 and DAQA. We provide a detailed analysis of the results for the different question types. We release the code to produce CLEAR2 as well as NAAQA to foster research in this newly emerging machine learning task.
Index Terms:
Acoustic question answering, acoustic reasoning, CNN, auditory scene analysis, CLEAR, Coordconv, Coordinate maps
1 Introduction
Question answering (QA) tasks are examples of constrained and limited scenarios for research in reasoning. The agent’s task in QA is to answer questions based on context. Text-based QA uses text corpora as context [undef, undefa, undefb, undefc, undefd, undefe]. In visual question answering (VQA) the questions are related to a scene depicted in still images [undeff, undefg, undefh, undefi, undefj, undefk, undefl]. Finally, video question answering attempts to use both the visual and acoustic information in video material as context [undefm, undefn, undefo, undefp, undefq, undefr]. The use of the acoustic channel is usually limited to linguistic information that is expressed in text form, either with manual transcriptions (e.g. subtitles) or by automatic speech recognition [undefs].
In most studies, reasoning is supported by spatial and symbolic representations in the visual domain [undeft, undefu]. However, reasoning and logic relationships can also be studied via representations of sounds [undefv]. Including the auditory modality in studies on reasoning is of particular interest for research in artificial intelligence, but also has implications in real world applications [undefw]. In [undefx], audio was used in combination with video and depth information to recognize human activities. It was shown that sound can be more discriminative than the corresponding visual cues. As an example, imagine using an espresso machine. Besides possibly a display, all information about the different phases of producing coffee, from grinding the beans, to pressing the powder into the holder and brewing the coffee with high pressure hot water are conveyed by the sounds. Detection of abnormalities in machinery where the moving parts are hidden, or the detection of threatening or hazardous events are other examples of the importance of the audio information for cognitive systems.
The audio modality provides important information that can be leveraged in the context of QA reasoning. Audio allows QA systems to answer relevant questions more accurately, or even to answer questions that are not approachable from the visual domain alone. In [undefy], we introduced the AQA task and proposed a new database (CLEAR) to promote research in AQA. The agent’s goal, in the proposed task, was to answer questions related to acoustic scenes composed by a sequence of elementary musical sounds. The questions foster reasoning on the properties of the elementary sounds and their relative and absolute position in the scene. To build CLEAR, we were inspired by the work of Johnson et al. [undeff] for VQA. Similarly, we tested an architecture built for VQA and based on FiLM layers [undefz] on the newly proposed AQA task. Next, [undefaa] [undefaa] proposed to extend the questions to more acoustically realistic situations by developing a new database called DAQA. To evaluate the results, they proposed the MALiMo network which relies on several FiLM layers.
The works cited above use neural network architectures that are largely inspired by image processing research. However, the structure of acoustic data is fundamentally different from that of visual data. This is illustrated for example in [undefab] where two standard data sets in computer vision (MNIST) and speech technology (Google Speech Commands) are compared via T-SNE [undefac]. A legitimate question is whether it is possible to obtain better results (in terms of accuracy and network complexity) by adapting the first layers of the architectures to take into account intrinsic characteristics of acoustic signals. Even within the AQA domain, the properties of acoustic data may vary significantly depending on the nature of the auditory scenes (e.g. CLEAR vs DAQA). It is, therefore interesting to evaluate the impact of the dataset on system performance.
To answer the above questions, we present a study that evaluates the impact of audio pre-processing, of acoustic feature extraction and of dataset characteristics on the performance neural architectures for AQA. When considering performance, we focus both on accuracy and complexity of the models. The main contributions can be summarized as follows:
  • We introduce CLEAR2 a more challenging version of the CLEAR dataset, which comprises scenes of variable duration and different elementary sounds for the training and test sets,
  • we regenerate a version of the DAQA dataset (DAQA)
  • we evaluate different ways of using a VQA architecture on AQA inputs,
  • we study the effect of time and frequency coordinate maps for acoustic data at different levels in the architecture,
  • we introduce new feature extraction modules motivated by the properties of acoustic data,
  • based on the above points, we propose a highly optimized FiLM-based architecture (NAAQA) that outperforms models that are 5 to 10 times more complex in terms of number of parameters,
  • we test the addition of a MALiMo module (from [undefaa]) into our NAAQA architecture,
  • we compare the performance NAAQA and NAAQA+MALiMo on CLEAR2 and DAQA’ to test generality of the methods,
  • we provide a detailed analysis of our results based on question type to improve interpretability.
The rest of the paper is organized as follows: Section 2 reports on recent related work, Section 3 describes both our CLEAR2 dataset and the DAQA dataset, Section 4 presents the QA models we have tested, Section 5 gives details on the experimental settings, Section 6 presents and discusses the results and, finally, Section 7 concludes the paper. Some extra information is reported in the supplementary material, and referred to in the paper through the corresponding section.
Figure 1: Overview of the CLEAR dataset generation process. Highlighted in red: 10 randomly sampled sounds from the elementary sounds bank, are assembled to create an acoustic scene. The attributes of each elementary sound are depicted in blue. The question template (orange) and the elementary sounds attributes are combined to instantiate a question. The answer is generated by applying each steps of the question functional program (purple) on the acoustic scene definition (blue). The impact of the reverberations can be seen in the changes of the signals envelops.
2 Related Work
This section presents previous research in QA systems and the needs for databases in AQA research.
2.1 Text-Based Question Answering
The question answering task was introduced as part of the Text Retrieval Conference [undef]. In text-based question answering, both the questions and the context are expressed in text form. Answering these questions can often be approached as a pattern matching problem in the sense that the information can be retrieved almost verbatim in the text (e.g. [undefb, undefc, undefe, undefd]).
2.2 Visual Question Answering (VQA)
Visual Question Answering aims to answer questions based on a visual scene. Several VQA datasets are available to the scientific community [undeff, undefad, undefae, undefaf, undefag, undefah, undefh, undefi, undefai, undefg]. However, designing an unbiased dataset is non-trivial. [undefj] [undefj] observed that the type of questions has a strong impact on the results of neural network based systems which motivated research to reduce the bias in VQA datasets [undefaj, undefak, undefal, undefk, undefl, undeff]. Gathering good labeled data is also non-trivial which induced [undefk] and [undefl] [undefk, undefl] to constrain their work to yes/no questions. To alleviate this problem, [undeff] [undeff] proposed the use of synthetic data for both questions and visual scenes. The resulting CLEVR dataset has been extensively used to evaluate neural networks for VQA applications [undefz, undefam, undefan, undefao, undefap, undefaq] which helped foster research on VQA. To create visual scenes, the authors automated a 3D modelling software. This allows for an unlimited supply of labeled data eliminating the time and effort needed for manual annotations. For the questions, they first manually designed semantic representations for each type of question. These representations describe the reasoning steps needed to answer a question (i.e. “find all cubes $|$ that are red $|$ and metallic”). The semantic representations are then instantiated based on the visual scene composition thus creating a question and an answer for a given scene. This gives complete control over the labelling process.
2.3 Dababases for AQA
As in VQA, using generated data in the design of AQA datasets has substantial advantages. Data can be automatically annotated which saves time and complexity. The number of training examples that can be generated is only limited by the available computational resources. Controlling the generation process gives a complete understanding of the properties and relations of the objects in a scene. This understanding can be leveraged to reduce bias in the dataset and to generate complex questions and their corresponding answers. The CLEAR dataset  [undefy] has been initially generated using semi-synthetic data. The elementary sounds were recordings of musical notes played by various instruments and players. However, the auditory scenes obtained by combining the elementary sounds had fixed duration. Moreover, the same elementary sounds were used to generate the test and training scenes. The DAQA dataset [undefaa] comprises more complex and less stationary elementary natural sounds coming for example from aircrafts, cars, doors, human speaking, bird singing, dog barking, etc. Although more complex and varied than CLEAR, the evaluation also uses the same elementary sound recordings for training and testing.
In this paper, we propose a more challenging version of CLEAR which uses different elementary sound recordings for the training and test sets and generates variable duration auditory scenes.
2.4 Convolutional neural network on Audio
Convolutional neural networks (CNN) have dominated the visual domain in recent years. More recently, they have also been applied to a number of problems in the acoustic domains such as acoustic scene classification [undefar, undefas, undefat, undefau], music genre classification [undefav, undefat], instrument classification [undefaw, undefax], sound event classification and localization [undefay] and speech recognition [undefat]. Some authors [undefav, undefaw, undefax, undefau] use intermediate representations such as STFT [undefaz], MFCC [undefaaa] or CQT [undefaab] spectrograms while others work directly with the raw audio signal [undefat, undefar].
Square convolutional and pooling kernels are often used to solve visual task such as VQA, visual scene classification and object recognition [undefaac, undefaad, undefaae, undefaaf]. [undefas, undefaag, undefaah] [undefas, undefaag, undefaah] have successfully used visually motivated CNN with square filters to solve audio related tasks. Time-frequency representations of audio signals are however structured very differently than visual representations. [undefaai] [undefaai] explore the performance of different structures of convolutive kernels when working with music signals classification. They propose the use of 1D convolution kernels to capture time-specific or frequency-specific features. They demonstrate that similar accuracy can be reached using a combination of 1D convolutions instead of 2D convolutions by combining 1D time and 1D frequency filters while using much fewer parameters. They also explore rectangular kernels which capture both time and frequency features at different scales. The impact of such strategies for music classification is still an open question in the context of auditory scene analysis.
Coordinate maps initially proposed in [undefaaj] by [undefaaj] have proven successful for processing visual data in the context of VQA. The method consists in augmenting the visual input with matrices containing numbers in the range -1 to 1 which vary either in the $x$ or in the $y$-dimension. With MALiMo [undefaa], the same strategy is used to indicate the simultaneous relative positions of features in frequency and time. [undefaak] [undefaak] proposed Frequency-Aware convolutions which are equivalent to concatenating coordinate maps only in the frequency axis. The effectiveness of coordinate maps on the time dimension for audio signals have not been studied to the best of our knowledge.
In this study we first evaluate the performance of a network initially designed for the VQA task (Visual FiLM) [undefz] on the AQA task. Then we introduce the NAAQA architecture to leverage specific properties of acoustic inputs. For this architecture, we analyze the influence of separate time and frequency coordinate maps. We then study the impact of adding a MALiMo block into our architecture. Finally, we evaluate our model on the DAQA dataset.
3 Data
We use two very different datasets in our experiments in order to study the effect of the AQA task characteristics on the model performance. The first set, that is also a contribution of this paper, comprises musical sounds (CLEAR2); the second includes short environmental sounds (DAQA).
3.1 CLEAR2
CLEAR2 is an updated version of CLEAR [undefy]. A graphical overview of the generation process is depicted in Figure 1. Each record in the dataset is a unique combination of a scene, a question and an answer.
Question type Example Possible Answers #
Note What is the note played by the flute that is after the loud bright D note? A, A#, B, C, C#, D, D#, E, F, F#, G, G# 12
Instrument What instrument plays a dark quiet sound in the end of the scene? bass, cello, clarinet, flute, trumpet, violin 5
Brightness What is the brightness of the first clarinet sound? bright, dark 2
Loudness What is the loudness of the violin playing after the third trumpet? quiet, loud 2
Absolute Position What is the position of the A# note playing before the bright B note? $\Big\}$ first, second … fifteenth 15
Relative Position Among the trumpet sounds which one is a F?
Global Position In what part of the scene is the clarinet playing a loud G note ? beginning, middle, end (of the scene) 3
Counting How many other sounds have the same brightness as the third violin? $\Big\}$0, 1 … 15 16
Counting Instruments How many different instruments are playing before the second trumpet?
[…]
Counting comparison Is there an equal number of loud cello sounds and quiet clarinet sounds?
Total 57
TABLE I: Types of questions with examples and possible answers. The variable parts of each question is emphasized in bold italics. The number of possible answer per question type is reported in the last column. Certain questions have the same possible answers, the meaning of which depends on the type of question.
To build acoustic scenes, we prepared a bank of elementary sounds composed of real musical instrument recordings extracted from the Good-Sounds [undefaal] dataset [^1] Each elementary sound in a scene is characterized by an n-tuple: [Instrument, Brightness, Loudness, Musical Note, Duration, Absolute position in scene, Relative position in scene, Global position]. The Brightness property is computed by using the timbralmodels [undefaam] library. A threshold is used to define the label of the sound (Dark or Bright). The Loudness labels are assigned based on the perceptual loudness as defined by the ITU-R BS.1770-4 international normalization standard [undefaan]. Again, a threshold is used to determine if the sound is Quiet or Loud. All attribute values are listed in Table I as possible answers to the questions explained below.. Differently from CLEAR, in CLEAR2 we make sure that the recordings (players, instruments, microphones) of the elementary sounds are different for the training and test sets. For the training set, the bank comprises 135 unique recordings (compared to 56 in CLEAR) sampled at 48KHz including 6 different instruments (bass, cello, clarinet, flute, trumpet and violin), 12 notes (chromatic scale) and 3 octaves. A different set of 135 recordings of the same instruments recorded using different microphones and players is used to create the test set. The acoustic scenes are built by concatenating between 5 to 15 randomly chosen sounds from the elementary sound bank into a sequence (as opposed to CLEAR which comprised fixed duration scenes). Silence segments of random duration are added in-between elementary sounds. The acoustic scenes are then corrupted by filtering to simulate room reverberation and by adding a white uncorrelated uniform noise. Both the amount of noise and reverberation vary from scene to scene with the goal of reducing overfitting.
For each scene, a number of questions is generated using CLEVR-like [undeff] templates. A template defines the reasoning steps required to answer a question based on the composition of the scene (i.e. “find all instances of violin $|$ that plays before trumpet $|$ that is the loudest”). 942 templates where designed for this AQA task. Not all template instantiations results in a valid question. The generated questions are filtered to remove ill posed questions similarly to [undeff]. Table I shows examples of questions with their answers.
The a priori probability of answering correctly with no information about the question or the scene, and assuming a uniform distribution of classes, is $\frac{1}{57}=1.75\%$. These probabilities are higher, on average, if we introduce information about the question. For example, if we know that the question is of the type Exist or Counting comparison, there are only two possible answers (yes or no) and the probability of answering correctly by chance is $0.5$. The majority class accuracy (always answering the most common answer: Yes) is 7.5%. Statistics on the CLEAR2 dataset are presented in Table II.
The generation process was built with extensibility in mind. Different versions of the dataset with fewer or more objects per scene can be generated by using different parameters for the generation script. It is also possible to modify the elementary sounds bank to generate datasets for AQA in other domains, speech or environmental sounds, for example. The code for generating the dataset is available on GitHub [^2] The link will be made available upon publication.
Datasets global statistics
Dataset CLEAR2 DAQA # of questions 200 000 599 441 # of scenes 50 000 100 000 # of answers 57 52 # of elementary sounds 135 ( + 135 for test) 358 # of types of question 11 5 # of unique vocabulary words 91 158
CLEAR2 Dataset detailed statistics
Mean Min Max # of sounds per scene 10 5 15 Elementary sound duration 0.85s 0.69s 1.11s Scene duration 10.69s 4.46s 17.82s # of words per question 17 6 28 # of unique words per question 12 5 19
DAQA’ Dataset detailed statistics
Mean Min Max # of sounds per scene 9 5 12 Elementary sound duration 9.35s 0.6s 20s Scene duration 1min 19s 9s 3min 4s # of words per question 13 5 27 # of unique words per question 11 5 22
TABLE II: Datasets statistics
3.2 Reproducing DAQA
We were not able to fully recreate the DAQA dataset because it relies on some AudioSet [undefaao] youtube videos that have since been deleted. We were able to retrieve 358 sounds out of the 400 sounds that were used to generate the original dataset. We used these sounds to generate the dataset. Changing the number of elementary sounds also impacts the whole generation process. This dataset is therefore different from the original DAQA and will be referred to as DAQA from now on. Our results are therefore not fully comparable to the ones reported in [undefaa]. A list of all the missing sounds is available in Supplementary G.
3.3 Comparing CLEAR2 and DAQA
Table II report statistics on both CLEAR2 and DAQA. The major difference between both datasets is the type of elementary sounds used to generate the acoustic scenes, that is, sustained musical notes (CLEAR2) versus possibily transient environmental sounds (DAQA).
Acoustic scenes in DAQA are much longer on average than the ones in CLEAR2. This results in much bigger input spectrograms, and, in turns, much higher computational requirements and longer training time. Besides computational costs, the reason for producing shorter acoustic scenes in CLEAR2 was to produce a task that could be solved by a human. It would be unreasonable to ask, e.g., if the “flute plays before the sixth violin” in a 2 minute audio clip.
Finally, the original DAQA [undefaa] and, consequently, our reconstruction (DAQA) suffer from the same problem as the original CLEAR. The same elementary sounds are used in the training and test scenes. Although scenes are still different between training and test set, this may cause the models to “remember” the elementary sounds rather than extracting their properties. In CLEAR2, this problem was mitigated by using different elementary sounds for the training and test set.
4 Method
We first describe the original Visual FiLM architecture [undefz] that we use as baseline model, then the proposed modifications that lead to our NAAQA architecture and, finally, NAAQA with a MALiMo module.
4.1 Baseline model: Visual FiLM
Both the proposed NAAQA and Visual FiLM [undefz] share an overall common architecture which is depicted in Figure 2. Visual FiLM, that we will use as baseline model, is inspired by Conditional Batch Normalization architectures [undefaap] and achieved state of the art results on the CLEVR VQA task [undeff]. The network takes a visual scene and a text-based question as inputs and predicts an answer to the question for the given scene. The text-processing module uses $G$ unidirectional gated recurrent units (GRUs) to extract context from the text input (yellow area in Figure 2). The visual scene is processed by the convolutional module (blue area in the figure). The first step of this module is feature extraction (orange box), performed by a Resnet101 model [undefaae] pre-trained on ImageNet [undefaaq]. The extracted features are processed by a convolutional layer with batch normalization [undefaap] and ReLU [undefaar] activation followed by $J$ Resblocks illustrated in details in the red area in the figure. Unless otherwise specified, batch normalization and ReLU activation functions are applied to all convolutional layers. Each Resblock $j$ comprises convolutional layers with $M$ filters that are linearly modulated by FiLM layers through the two $M\times 1$ vectors $\boldsymbol{\beta}_{j}$ (additive) and $\boldsymbol{\gamma}_{j}$ (multiplicative). This modulation emphasizes the most important feature maps and inhibits the irrelevant maps given the context of the question. $\boldsymbol{\beta}_{j}$ and $\boldsymbol{\gamma}_{j}$ are learned via fully connected layers using the text embeddings extracted by the text processing module as inputs (purple area in the figure). The affine transformation in the batch normalization before the FiLM layer is deactivated. The FiLM layer applies its own affine transformation using the learned $\boldsymbol{\beta}_{j}$ and $\boldsymbol{\gamma}_{j}$ to modulate features. Several Resblocks can be stacked to increase the depth of the model, as illustrated in Figure 2. Finally, the classifier module is composed of a $1\times 1$ convolutional layer [undefaas] with $C$ filters followed by max pooling and a fully connected layer with $H$ hidden units and an output size $O$ equal to the number of possible answers (Gray in Figure 2). A softmax layer predicts the probabilities of the answers. In order to use the Visual FiLM as a baseline for our experiments, we extract a 2D spectro-temporal representation of the acoustic scenes as depicted at the bottom of Figure 2. The Resnet101 pre-trained extractor expects a 3 channels visual input but the spectro-temporal representation comprises only 1 channel. To work around this constraint, the spectro-temporal information is simply repeated 3 times thus creating a 3 channels input (only when using Resnet101 as feature extractor). This modified spectro-temporal representation is then fed to the model as if it was an image which is the simplest way to adapt the unmodified visual architecture to acoustic data. We call this architecture Visual FiLM Resnet101.
Figure 2: Common Architecture. Two inputs: a spectro-temporal representation of an acoustic scene and a textual question. The spectro-temporal representation goes through a feature extractor (Parallel and Interleaved feature extractor detailed in Section 4.2.1 for NAAQA and Resnet101 pretrained on ImageNet for Visual FiLM) and then a serie of $J$ Resblocks that are linearly modulated by $\boldsymbol{\beta}_{j}$ and $\boldsymbol{\gamma}_{j}$ (learned from the question input) via FiLM layers. Coordinate maps are inserted before convolution blocks with a pink border. The output is a probability distribution of the possible answers.
(a) Parallel feature extraction. The input spectrogram is processed by 2 parallel pipelines. The first pipeline (in green) captures frequency features using a serie of $K$ 1D convolutions with $N_{k}$ filters and a stride of $2\times 1$. Since the stride is larger than $1\times 1$, each convolution downsample the frequency axis. The $1\times 2$ maxpooling then downsamples the time axis. The second pipeline (in yellow) captures time features using the same structure with transposed filter size. Features from both pipelines are concatenated and fused using a $1\times 1$ convolution with $P$ filters to create.
(b) Interleaved feature extraction. 1D time (in yellow) and frequency (in green) convolutions are applied alternately on the input spectrogram building a time-frequency representation after each block. The order of the convolution in each block can be reversed. The extractor is composed of $K$ blocks where each convolution has $N_{k}$ filters followed by a $1\times 1$ convolution with $P$ filters.
Figure 3: Acoustic feature extraction
4.2 The proposed NAAQA architecture
In order to define the NAAQA architecture, we performed a number of modifications to the baseline architecture that will be described in the following sections.
4.2.1 Feature Extraction
As in Visual FiLM, the first step in the NAAQA model is feature extraction (orange box in Figure 2). The most obvious adaptation of Visual FiLM to acoustic data is to retrain the feature extraction module on the scenes from CLEAR2. In order to do this, we used three 2D convolutional layers, with $3\times 3$ kernels, stride $2\times 2$ and $N_{1}$, $N_{2}$, and $N_{3}$ filters respectively followed by a $1\times 1$ convolutions with $N_{4}$ filters. We will refer to this model as NAAQA 2D Conv.
However, because acoustic signals present unique properties, we introduce two feature extraction modules that are specifically tailored to modeling sounds: the Parallel feature extractor (Figure 3a) processes time and frequency features independently in parallel pipelines; the Interleaved feature extractor (Figure 3b) captures time and frequency features in a single convolutional pipeline. In both cases, the feature extractor is trained end-to-end with the rest of the network and uses a combination of 1D convolutional filters to process a 2D spectro-temporal representation. The 1D filters process the time and frequency axis independently as opposed to the 2D filters typically used in image processing.
The design of the Parallel feature extractor (Figure 3(a)) is inspired by the work of [undefaai] [undefaai] where 1D filters are used to capture time and frequency features separately. While [undefaai] time-frequency model includes only 1 time and 1 frequency convolution which are then concatenated together, our extractor stacks multiple 1D convolutions in two parallel pipelines. The time and frequency features are only fused at the end of both pipelines. This yields more complex features. The frequency pipeline (green in the figure) comprises a serie of $K$ frequency blocks. Each block is composed of a 1D convolution with $N_{K}$ $3\times 1$ kernels and $2\times 1$ strides followed by a $1\times 2$ maxpooling. With a stride larger than $1\times 1$, the convolution operation downsamples the frequency axis and the pooling operation downsamples the time axis. This downsampling strategy allows features in both parallel pipelines to be of the same dimensions. The time pipeline (yellow in the figure) is the same as the frequency pipeline except that the convolutional kernel operates along the time dimension and the pooling along the frequency dimension. The convolution kernel is $1\times 3$ and the pooling kernel is $2\times 1$. The activation maps of both pipelines are concatenated channel-wise and a representation combining both the time and frequency features is created using a $1\times 1$ convolution [undefaas] with $P$ filters and a stride of one. The feature maps dimensionality is either compressed or expanded depending on the number of filters $P$ in the $1\times 1$ convolution. We call the corresponding model NAAQA Parallel. The $1\times 1$ convolution can also be removed thus leaving it up to the next $3\times 3$ convolution to fuse the time and frequency features.
The Interleaved feature extractor (Figure 3(b)) processes the input spectrogram in a single pipeline composed of a serie of $K$ interleaved blocks (purple in the figure). Each block comprises a $1\times 3$ time convolution with $N_{K}$ filters and stride $1\times 2$ followed by a $3\times 1$ frequency convolution with $N_{K}$ filters and stride $2\times 1$. A $1\times 1$ convolution with $P$ filters processes the output of the last block to either compress or expand its dimensionality. We call the corresponding model NAAQA Interleaved Time.
As an alternative configuration, the order of the convolution operation in each block can be reversed so that it first operates along the frequency axis and then the time axis. The model is called NAAQA Interleaved Freq, in this case. Compared to the Parallel feature extractor, time-frequency representations are created after each block instead of only at the end of the pipeline.
For all extractors, the convolutions in the first block comprise $N_{1}$ convolutional filters and the number of filters is doubled after each block ($N_{i}$ = $2N_{i-1}$). More blocks (higher $K$) results in more down sampling of the feature maps shape which brings down the computational cost of the model.
4.2.2 Coordinate maps for acoustic inputs
The Visual FiLM model concatenates coordinate maps (CoordConv [undefaaj]) to the input of convolutional layers (pink border boxes in Figure 2) when tackling the VQA task. In the visual domain both axis of an image encode spatial information. Coordinate maps have, therefore, the same meaning in the $x$ or $y$-axis.
In spectro-temporal representations for audio, however, the $y$-axis corresponds to frequency and the $x$-axis to time. We, therefore, call the maps frequency and time coordinate map, respectively. All spectro-temporal representations in CLEAR2 have the same range for the frequency axis but the range for the time axis varies depending on the duration of the acoustic scenes. We hypothesize that time coordinate maps might have a stronger impact on performance because they provide a relative time scale that the model can use to enhance its temporal localization capabilities.
4.2.3 Complexity Optimization
We performed optimization of the most important hyper-parameters in the NAAQA architecture with the goal of reducing model complexity. These include number of GRU text-processing units $G$; the number of Resblock $J$ that dictates the number of FiLM layers and, therefore, the number of modulation coefficients to compute; the number of convolutional filters $M$ in each block; the number of filters $C$ and the number of hidden units $H$ in the classifier module. We refer to the resulting model by prepending Optimized to the model name.
4.2.4 NAAQA with a MALiMo module
In MALiMo [undefaa], [undefaa] add a second set of FiLM layers that acts has an auxiliary controller. The controller uses the extracted acoustic features to further modulate the intermediate Resblocks. To evaluate the impact of MALiMo on CLEAR2, a MALiMo module was added to NAAQA. We refer to this configuration by appending MALiMo ctrl to the names introduced above. In our implementation of the module we replaced LSTMs with GRUs and adapted the inputs to the acoustic features that we study.
5 Experiments
We perform experiments to compare the effect on performance of our modification to the baseline model. Most of our experiments are performed on the proposed CLEAR2 data. We first investigate the performance of different feature extraction methods and compare them to the Visual FiLM Resnet101 feature extractor. Then, we show the effect of time and frequency coordinate maps at different levels of the model. Moreover, we perform an hyper-parameters ablation study to reduce the complexity of the model. We finally test the addition of a MALiMo module to our model. In order to demonstrate the generality of our results, we compare the performance of our model on the CLEAR2 and DAQA datasets.
5.1 Acoustic Pre-processing
The raw acoustic signal (sampled at 48 kHz for CLEAR2 and 16kHz for DAQA) is processed to create a 2D time-frequency representation using Mel scale [undefaat] spectrograms. After some experimentation we decided to use 64 Mel coefficients for both CLEAR2 and DAQA computed over samples weighted by a Hanning window. For CLEAR2 the window size was of 512 samples ($\sim$10.6 msec) whereas for DAQA we used windows of 400 samples ($\sim$25ms) as in [undefaa]. The time shift between consecutive windows (stride) was also optimized depending on the properties of the audio data. For CLEAR2 we found the best results for a time shift of 2048 samples ($\sim$42.7 msec). This is possible because of the properties of the sustained notes in the data set which vary slowly in time. Using such a long time shift allowed us to reduce more than ten folds the computational costs in our experiments. Because the DAQA dataset contains sounds that are more quickly varying, the same optimization is not possible. We tested with time shifts of 160 samples (10ms) and 1600 samples (100ms) observing a 5% drop in accuracy for the longer stride. All results based on CLEAR2 are reported with long window shift (long stride), with the exception of the comparison between short and long stride on both CLEAR2 and DAQA in Supplementary C.
Because the duration of the scenes in CLEAR2 is not constant, we pad the spectrogram along the time axis with zeros so that all spectrograms have the same dimensions ($1\times 64\times 418$) which corresponds to a maximum length of $\sim$17.9 sec. The power spectrum is normalized to the mean and standard deviation of the training data with the goal of speeding up convergence [undefaau].
5.2 Experimental conditions
Unless specified otherwise, the models presented in subsequent sections are trained on the CLEAR2 dataset which comprises 50 000 scenes and 4 questions per scene for a total of 200 000 records from which 140 000 (70%) are used for training, 30 000 (15%) for validation and 30 000 (15%) for test. The test set is generated using a different set of elementary sound which ensure that the network can not memorize the elementary sounds and therefore acts has a better benchmark for generalization. Optimisation technique and other training settings are further described in Supplementary A. The results are reported in terms of accuracy, that is the percentage of correct answers on the total. Since initialization of deep architectures has a profound impact on training convergence, we developed a python library torch-reproducible-block [^3] The link will be made available upon publication to control the model initial conditions and design reproducible experiments. To ensure the robustness of the results, each model is trained 5 times with 5 different random seeds.
5.3 Initial model configuration
The initial configuration for the proposed model comprises $G=4096$ GRU units, $J=4$ Resblocks with $M=128$ filters each and a classifier composed of a $1\times 1$ convolution with $C=512$ filters and $H=1024$ hidden units in the fully connected layer. This configuration includes both time and frequency coordinate maps in each location highlighted in pink in Figure 2.
Number of Overall Accuracy by question type (%)
Configuration Parameters Acc. Instrument Note Brightness Loudness Exist Abs. Pos. Glob. Pos. Rel. Pos. Count Count Comp. Count Inst.
Baselines
Random Answer 1.75 1.75 1.75 1.75 1.75 1.75 1.75 1.75 1.75 1.75 1.75 1.75
Most common Answer (Yes) 7.3 0 0 0 0 55.62 0 0 0 0 47.27 0
Visual FiLM Resnet101 6.71 M 62.3 ±0.78 61.9 ±0.85 37.8 ±1.30 83.9 ±0.58 81.9 ±0.61 73.2 ±0.70 51.7 ±1.40 73.1 ±2.29 48.4 ±1.82 41.5 ±0.42 59.8 ±0.57 39.9 ±3.03
NAAQA
NAAQA 2D Conv 5.61 M 77.6 ±0.72 80.6 ±0.66 73.4 ±1.28 90.5 ±1.44 86.5 ±0.76 81.3 ±0.57 74.7 ±2.50 87.5 ±0.42 54.0 ±1.43 53.8 ±1.01 60.8 ±1.34 49.0 ±3.53
NAAQA Interleaved Freq 5.61 M 67.2 ±0.98 63.0 ±1.93 48.0 ±2.14 84.5 ±1.07 81.1 ±0.52 72.1 ±1.24 65.1 ±1.04 80.1 ±0.66 49.3 ±1.59 42.5 ±1.51 56.6 ±2.45 46.5 ±1.57
NAAQA Interleaved Time 5.61 M 78.0 ±0.51 81.8 ±1.06 70.6 ±1.24 90.9 ±0.78 87.9 ±0.94 81.6 ±0.66 75.9 ±1.08 87.4 ±0.25 60.0 ±4.07 53.6 ±0.63 61.2 ±0.73 50.0 ±2.23
NAAQA Parallel 5.61 M 78.5 ±0.45 80.4 ±0.83 72.6 ±1.19 91.2 ±0.74 87.1 ±0.28 80.4 ±0.16 78.7 ±1.26 88.8 ±0.53 55.4 ±1.63 52.6 ±0.43 59.8 ±2.13 50.2 ±3.49
Optimized NAAQA Parallel 1.68 M 79.5 ±0.05 81.7 ±0.20 74.2 ±0.58 91.9 ±0.17 87.4 ±0.31 81.2 ±0.10 79.3 ±0.77 90.0 ±0.15 58.0 ±1.32 53.8 ±1.40 60.6 ±0.89 50.4 ±1.10
NAAQA + MALiMo
Visual FiLM Resnet101 + MALiMo ctrl 6.82 M 63.6 ±1.20 61.8 ±1.40 36.3 ±1.41 83.5 ±0.86 82.3 ±0.59 73.4 ±1.06 57.1 ±3.58 76.1 ±1.94 47.4 ±2.98 41.6 ±0.80 56.7 ±4.23 43.2 ±3.64
NAAQA 2D Conv + MALiMo ctrl 6.71 M 77.1 ±1.12 79.2 ±1.01 71.3 ±1.54 90.5 ±0.81 86.1 ±0.73 80.3 ±0.77 75.7 ±2.08 87.8 ±0.74 52.7 ±1.94 51.9 ±2.73 59.0 ±3.00 50.2 ±2.48
NAAQA Interleaved Freq + MALiMo ctrl 6.71 M 64.8 ±3.91 55.3 ±8.33 43.1 ±8.77 83.2 ±1.33 80.0 ±1.75 70.3 ±2.74 63.6 ±4.72 78.7 ±3.27 45.8 ±3.53 41.4 ±1.96 55.1 ±2.75 48.7 ±2.05
NAAQA Interleaved Time + MALiMo ctrl 6.71 M 77.1 ±0.63 79.9 ±0.86 70.8 ±1.10 90.5 ±1.00 86.7 ±0.89 80.3 ±0.34 74.5 ±1.06 87.6 ±0.35 55.2 ±2.39 53.8 ±0.43 59.9 ±2.32 50.0 ±0.93
NAAQA Parallel + MALiMo ctrl 6.71 M 77.3 ±0.93 78.2 ±1.50 71.3 ±1.00 89.9 ±0.66 85.8 ±0.67 79.7 ±0.45 77.7 ±2.02 87.6 ±0.53 53.6 ±1.84 51.9 ±1.82 59.0 ±2.29 48.1 ±1.65
Optimized NAAQA Parallel + MALiMo ctrl 2.78 M 78.2 ±0.06 80.8 ±0.41 72.4 ±0.33 89.6 ±0.62 86.2 ±0.04 79.7 ±0.10 79.3 ±0.22 88.4 ±0.36 54.0 ±0.33 52.2 ±0.70 58.8 ±0.45 45.6 ±0.00
TABLE III: Results on CLEAR2. The table shows number of parameters, average accuracy (%), and standard deviation over 5 repetitions of the training. Overall accuracy as well as question-kind dependent accuracy are reported. The different configurations are reported in the same order as they are presented in the paper. The most common answer is “Yes”.
Configuration # Parameters Train Acc. Val Acc. Test Acc. Trainig time
Optimized NAAQA 2D Conv 1.68 M 65.2 ±0.64 58.0 ±0.82 58.3 ±0.98 0 days 06:48:12
Optimized NAAQA Parallel 1.68 M 66.6 ±0.51 60.4 ±0.08 60.4 ±0.21 0 days 06:49:52
Optimized NAAQA 2D Conv + MALiMo ctrl 2.78 M 58.5 ±3.03 54.2 ±2.42 54.4 ±2.35 0 days 07:31:37
Optimized NAAQA Parallel + MALiMo ctrl 2.78 M 67.3 ±1.22 64.1 ±0.54 64.3 ±0.72 1 days 04:59:46
TABLE IV: Results on DAQA The table presents number of parameters, average training, validation and test accuracy (%) with standard deviation over 5 repetitions of the training as well as average training time. Results are reported for four configurations, with and without the MALiMo module in the same order as they are presented in the paper.
6 Results and Discussion
Main results on the CLEAR2 data set are reported in Table III. It reports the complexity of the models in terms of number of parameters, the overall accuracy and accuracy dependent on question’s type. First we report two theoretical baselines given by random chance and majority class answers. Then we report results based on the baseline model Visual FiLM Resnet101 using the initial configuration described in Section 5.3. This architecture achieves the lowest accuracy of all tested models with 62.3%. This suggests that Mel spectrograms have a very different structure than visual scenes and, as expected, the pre-learned knowledge gathered in a visual context does not transfer directly to the acoustic context.
6.1 NAAQA modifications
Unless specified otherwise, the initial configuration described in section 5.3 is used for all models in this section.
6.1.1 Feature Extraction
The first improvement to the baseline is given by introducing a specific audio feature extraction module based on 2D convolutions. The NAAQA 2D Conv model has slightly fewer parameters than Visual FiLM Resnet101 because of the simpler feature extraction module and a much higher overall accuracy of 77.6%.
We then tested two versions of the Interleaved feature extractor (Figure 3(b)). Results show that the order of the 1D convolutions in each block has a significant impact on performance. When the first 1D convolution in each block is done along the frequency axis (NAAQA Interleaved Freq), the network reaches an overall accuracy of 67.2%. It performs especially poorly with questions related to count (42.5%), count instruments (46.5%) and notes (48.0%). The performance on position questions is also the lowest among all the extractors. When the order of the convolution is reversed (NAAQA Interleaved Time), useful information is better captured and the network reaches 78.0% of overall accuracy. A possible explanation relates to the nature of the sounds in the CLEAR2 dataset which mainly consists of sustained musical notes. The time dimension at short scales does not contain much information that helps identifying the individual sounds. At larger scales, however, the time axis contains information relative to the scene as a whole which is exploited by higher level layers (Resblocks) to model the connections between different sounds. Because its stride is greater than 1x1, each 1D convolution downsamples the axis on which it is applied. When the first convolution is a frequency convolution, the frequency axis of the resulting features is downsampled which reduces the useful information that can be captured by the time convolution that follows.
The Parallel feature extractor (NAAQA Parallel, Figure 3(a)) reaches an overall accuracy of 78.5%. It performs well on all question’s types except relative position, count and count instrument. Refer to Section 6.2 for further analysis. These results show that building complex time and frequency feature separately and fusing them at a later stage is a good strategy to learn acoustic features for this task. This claim is further strengthen by the analysis of section 6.3.2.
Out of all extractors, NAAQA Parallel is the one that performs the best and constitutes the basis of NAAQA in all subsequent experiments.
6.1.2 Coordinate Maps
Coordinate maps Accuracy (%)
Extractor 1st Conv Resblocks Classifier Train val Test
Time 95.0 ±0.79 90.2 ±0.62 79.0 ±0.44
Time 95.1 ±0.90 90.0 ±0.28 79.0 ±0.43
Both 94.9 ±1.08 90.1 ±0.83 78.8 ±0.52
Both 95.1 ±1.23 90.0 ±0.46 78.7 ±0.70
Time 94.7 ±1.06 88.3 ±1.24 76.7 ±0.83
Both 94.1 ±1.47 88.4 ±1.06 76.6 ±0.36
Freq 84.0 ±2.17 73.5 ±1.40 64.8 ±1.51
85.2 ±0.71 72.9 ±1.36 63.8 ±0.70
Both 83.5 ±3.46 72.3 ±2.71 63.5 ±1.93
Freq 84.5 ±0.67 71.7 ±2.85 62.6 ±2.36
Freq 83.4 ±1.53 70.7 ±1.74 61.7 ±1.16
Time 81.9 ±0.88 70.1 ±1.99 61.7 ±0.76
Freq 79.7 ±3.70 67.1 ±3.36 59.3 ±2.20
TABLE V: Impact of the placement of Time and Frequency coordinate maps. All possible positions are illustrated by the pink border boxes in Figure 2. The value Both indicate that both Time and Frequency coordinate maps were inserted at the given position. The NAAQA Parallel is used with hyper-parameters from the initial configuration (defined in section 5.2. The rows are ordered by test accuracy.).
Coordinate maps can be inserted before any convolution operation (Figure 2). We therefore analyzed the impact of the placement of Time and Frequency coordinate maps at different depths in the network. All possible locations were evaluated via grid-search. For each location, we inserted either a Time coordinate map, a Frequency coordinate map or both. Results are detailed in Table V. Time coordinate maps have the biggest impact on performance, especially when inserted in the first convolution after the feature extractor or in the Resblocks. This could be because the fusion of the textual and acoustic features, and therefore most of the reasoning, is performed in the Resblocks. The network might be using the additional localization information to inform the modulation of the convolutional feature maps based on the context of the question. Surprisingly, the Frequency coordinate maps have a minimal impact on performance. We further compare the impact of Time versus Frequency coordinate maps in Supplementary B.
6.1.3 Complexity Optimization
Figure 4: Test accuracy by question type and the number of relation in the question for Optimized NAAQA Parallel. The overall accuracy for this configuration is 79.1%. The presence of before or after in a question constitutes a temporal relation. The accuracy is N/A for relative position and count compare since these types of question do no include relations. The hyper-parameters used are described at the end of Section 6.1.3
As described in Section 4.2.3, we optimized the most important hyper-parameters ($G,J,M$) in the NAAQA model in order to reduce its complexity. The baseline Visual FiLM Resnet101 configuration comprises 6.71 M parameters and achieves only 62.3%. NAAQA Parallel comprises 5.61 M parameters and performs significantly better with 78.5%. With this model as starting point, we performed an ablation study to find which hyper-parameters can be reduced without impacting accuracy. The Optimized NAAQA Parallel configuration is the best trade-off between model complexity and performance. It comprises 1.68 M parameters and achieves the best overall accuracy with 79.5%. The most notable complexity reduction comes from the reduction of the number of GRU units $G$. Reducing $G$ from $4096$ to $512$ increased accuracy while reducing the number of parameters by a factor of 3 (6.61 M vs 1.68 M). The Optimized NAAQA Parallel is composed of a Parallel extractor with $K=3$ blocks and $P=64$, $G=512$ GRU units, $J=4$ Resblocks with $M=128$ filters, a classifier module with $C=512$ filters and $H=1024$ units. Results for this configuration can be found in Table III and Figure 4. Further results related to the ablation study can be found in Supplementary H.
6.1.4 Adding a MALiMo controller
The bottom rows of Table III show results where the configurations described in previous sections are augmented with a MALiMo controller. Although the model complexity is significantly increased ($\sim$1M parameters), this addition does not bring any improvement in the model performance on CLEAR2. Almost all the tested configurations with a MALiMo controller perform slightly worse than the same configuration without the module, as can be seen in Table III. This may be again explained by the characteristics of the sounds in CLEAR2. A more in-depth discussion is given when we evaluate the models on DAQA.
6.2 Summary of Results on CLEAR2
NAAQA performs well on the CLEAR2 AQA task with 79.5% overall accuracy. It does however struggle with certain types of question as shown in Table III and Figure 4. When asked to count the number of sounds with specific attributes, NAAQA reaches only 53.8% accuracy. This limitation is more severe if the question is to count the different instruments playing in a given part of the scene (50.4%). It attains slightly higher accuracy when asked to compare the number of instances of acoustic objects (more, fewer or equal number) with specific attributes (60.6%). In contrast, the network can successfully recognize individual instruments in the scene (81.7%). This suggests, that the problem lies in the logical complexity of the question rather than in the pattern matching from the acoustic scene. As an example, the question (count instrument): ”How many different instruments are playing after the third cello playing a C# note?” requires to first identify the cello playing the C# note, then identify all acoustic objects that are playing after this sound, determine which instruments are of the same family and finally count the number of different families. The model struggles when it must focus on a large number of acoustic objects which explain the low accuracy for this type of question.
A similar argument could explain why models also have difficulties with questions related to the relative position of the instruments (58.0%). For example, to answer the question ”Among the flute sounds, which one plays an F note?”, the model must find all the flutes playing in the scene, determine which one play an F note, count the number of flute playing before and translate the count to a position [^4] This is only one possible strategy to answer the question. There may be other ways.. This also requires the network to focus on multiple objects.
Certain questions include temporal relations between sounds (before and after) as exemplified in Table I. Questions that include relations require focusing on several sounds to be answered. Figure 4 shows the accuracy for each question type depending on the presence of temporal relations. Questions that require the network to focus on a single acoustic object (brightness, loudness, instrument, note, global position and absolute position) benefit from the presence of a relation in the question. This might be explained by the fact that the question contains more information about the scene which helps to focus on the right acoustic object. However, the presence of relations in questions that already require the network to focus on multiple objects (exist, count and count comparison) is detrimental. This again supports the idea that having to focus on too many objects in the scene hinders the network performance.
6.3 Evaluation of DAQA
In order to compare our results to those of [undefaa] [undefaa], we evaluated our models on a version of the DAQA data set. As mentioned in Section 3.2, we were not able to reproduce the original DAQA dataset which mean that the results presented in this section are not fully comparable with [undefaa]. We report results for different configurations of NAAQA on our modified DAQA in Table IV.
6.3.1 NAAQA on DAQA
The models explored in this section are considerably smaller in terms of number of parameters than the ones explored in [undefaa]. The smallest model they evaluated had 5.49M parameters, the biggest model had 21.33M parameters and the best performing model had 13.20M parameters. The Optimized NAAQA 2D Conv model only has 1.68 M parameters and reaches an accuracy of 58.3% on DAQA. The Optimized NAAQA Parallel has the same number of parameters and performs slightly better with and accuracy of 60.4%. When we analyzed both of these models on CLEAR2 dataset in Section 6.1.1, we found a much smaller difference between the performance of the NAAQA 2D Conv and the NAAQA Parallel. This difference in the observations suggests that the parallel extractor is more useful in the context of complex acoustic sounds (DAQA) than with sustained musical note (CLEAR2).
Even though these results are not 100% comparable with [undefaa] because of the difference in the dataset composition, we want to emphasize that Optimized NAAQA Parallel reaches a somewhat similar accuracy than the smallest FiLM in [undefaa] (60.4% vs 64.3%) with significantly smaller number of parameters (1.68 M vs 5.49 M).
6.3.2 NAAQA with a MALiMo module on DAQA
In Section 6.1.4, we found that adding a MALiMo controller to our NAAQA models did not improve the accuracy on CLEAR2. Results in Table IV show that the MALiMo controller have a significant impact on performance when the model is evaluated on DAQA dataset. We see an increase of almost 4% when using Optimized NAAQA Parallel + MALiMo ctrl compared to Optimized NAAQA Parallel alone. These results are consistent with [undefaa] findings and with the hypothesis that MALiMo increases performance when working with complex sounds.
The Optimized NAAQA Parallel + MALiMo ctrl configuration performs about the same as the smallest MALiMo model evaluated in [undefaa] (64.3% vs 65.1%) with significantly less parameters (2.78 M vs 8.91 M).
7 Conclusions
Acoustic Question Answering (AQA) is a newly emerging task in the area of machine learning research. As performance is strongly dependent on the acoustical environments and types of questions, it is important to understand the relationship between the application and the chosen neural architecture. We propose a benchmark for AQA based on musical sounds (CLEAR2) and a neural architecture that is tailored to interpreting acoustic scenes (NAAQA). NAAQA introduces a number of modifications to a FiLM based architecture in order to optimize acoustic scenes analysis. These includes several strategies for neural feature extraction, an ablation study of the hyper-parameters and the optimization of coordinate maps for this task. We confirm that FiLM layers are very effective to modulate activation maps in the AQA application. We are able to optimize our NAAQA neural network so to obtain competitive results with a fraction of the model complexity. These results are confirmed on a different AQA task (DAQA) comprising more complex sounds with the addition of a MALiMo controller in the model. We release all code openly in the hope that these resources may foster increased research activity in solving the AQA task.
Acknowledgment
We would like to acknowledge the NVIDIA Corporation for donating a number of GPUs. Part of this research was funded or supported by the CHIST-ERA IGLU project, the CRSNG, the Michael-Smith scholarships the University of Sherbrooke and NTNU.
Supplementary Material
Supplementary A Detailed training settings
All models were trained for a maximum of 40 epochs or until the validation loss stops decreasing for 6 consecutive epochs. The Adam optimizer [undefaav] with a learning rate of $3\text{\times}{10}^{-4}$ was used to optimize the cross entropy loss. To reduce the likelihood of the model getting trapped in a local minima, the learning rate is automatically reduced by a 0.1 factor if the validation loss stops decreasing for 3 consecutive epochs. L2 regularization with a weight decay of $5\text{\times}{10}^{-6}$, dropout [undefaaw] with a drop probability of 0.25 and batch normalization [undefaap] are used to regularize the learning process. A batch size of 128 is used for all experiments. The following 5 random seeds were used to run all experiments : 876944 189369 682421 175326 427438
Supplementary B Time and Frequency coordinate maps impact
Accuracy by question type on the test set (%)
Coordinate Absolute Global Relative Count Count
Maps Overall Instrument Note Brightness Loudness Exist Position Position Position Count Comparison Instruments
Time only 79.0 ±0.43 80.8 ±1.22 73.0 ±0.84 90.8 ±0.52 86.8 ±0.49 80.2 ±0.75 80.2 ±0.89 89.0 ±0.60 53.1 ±0.90 52.1 ±0.70 61.1 ±0.56 49.3 ±0.47
Time & Freq 78.7 ±0.70 80.3 ±0.92 72.5 ±0.82 90.8 ±0.83 87.0 ±1.02 80.6 ±0.47 79.9 ±1.25 88.9 ±0.58 54.8 ±1.21 52.9 ±0.42 59.1 ±2.46 46.0 ±3.14
None 63.8 ±0.70 73.3 ±0.72 61.8 ±0.62 87.6 ±0.66 84.6 ±0.96 77.1 ±0.56 41.0 ±2.19 69.8 ±1.33 49.0 ±2.02 45.9 ±0.80 59.3 ±0.41 44.6 ±1.71
Freq only 61.7 ±1.16 72.7 ±0.88 60.4 ±0.89 87.4 ±0.62 83.4 ±0.41 75.6 ±0.71 36.8 ±3.80 67.2 ±1.58 47.6 ±1.86 44.2 ±1.58 57.0 ±2.80 43.5 ±4.09
TABLE VI: Comparison of Time and Frequency coordinate maps. Coordinate maps are inserted in the Resblocks. Test accuracy is reported by question type (Examples for each question type are available in Table I).
The impact of the kind of coordinate maps on each question types is presented in Table VI. To facilitate comparison, all coordinate maps are inserted at the same location (in the first convolution after the feature extractor). When the network does not include any coordinate maps it reaches only 63.8%. Inserting a Frequency coordinate map decreases accuracy to 61.7%. As hypothesized, Time coordinate maps significantly increase the overall accuracy with 79%. The accuracy increase for all question types is most notable with questions related to Absolute position where the accuracy goes from 41.0% to 80.2%. This is a similar result to [undefaaj] where coordinate maps were proved to help convolutional networks locate objects in visual scenes. In the context of a spectro-temporal representation, this ability translates to enhanced temporal localization which naturally helps to answer questions related to position. Questions related to Notes also receive a significant boost from Time coordinate maps going from 61.8% to 73%. The fact that Frequency coordinate maps decrease performance on questions related to Notes is rather surprising since notes are mostly defined by the frequency content at a given point in time. Furthermore, combining both Time and Frequency coordinate maps slightly reduces the overall performance when compared to using only Time coordinate maps which is in line with previous results.
Supplementary C Results showing Effects of stride size on both CLEAR2 and DAQA’
Model Nb Parameters Train acc Val acc Test acc Trainig time
On CLEAR2 dataset
Optimized NAAQA Parallel- Long stride 1.68 M 95.2 ±0.43 90.5 ±0.28 79.1 ±0.32 0 days 02:39:51
Optimized NAAQA Parallel- Short stride 1.68 M 93.1 ±0.48 89.5 ±0.38 77.9 ±0.68 0 days 07:00:01
On DAQA dataset
Optimized NAAQA Parallel- Short Stride 1.68 M 68.0 ±0.22 64.6 ±0.34 65.3 ±0.37 3 days 17:27:45
Optimized NAAQA Parallel- Long Stride 1.68 M 67.3 ±1.38 60.2 ±0.24 60.3 ±0.20 0 days 07:14:17
TABLE VII: Short stride and Long Stride comparison The optimized NAAQA on CLEAR2 and DAQA’. The rows are ordered by test accuracy.
Supplementary D Discussion on results with Resnet
Multiple feature extractors were evaluated in Section 6.1.1. One of them is the Resnet-101 pre-trained on ImageNet (Visual dataset). This extractor achieves the lowest of all tested extractors with 62.3%. However, it performs quite well with questions related to brightness and loudness with 83.9% and 81.9% respectively. One possible explanation is that both brightness and loudness are directly related to the amplitude of the “pixels” along the frequency axis which may be efficiently encoded by the Resnet101 model. The baseline does not perform as well with questions related to absolute position which is somewhat surprising since these questions correspond to a similar pattern matching problem on the time axis instead of the frequency axis. The baseline model also has difficulties with questions related to notes (37.8%). This might be explained by the fact that a note can be identified by its fundamental frequency and harmonics which can be far apart on the frequency axis. Visual models are trained to recognize localized features and therefore struggle to recognize notes. This is a feature that is typical of acoustic signals.
It does however perform better with global position questions. This is logical because this type of question refers to an approximate position (beginning, middle, end) which constrain the number of possible answers to 3 instead of 15.
Supplementary E Variability of the input size
Batch normalization has been proven to help regularize the training of neural networks [undefaap]. However, using batch normalization requires all inputs in a batch to have the same dimensions in order to compute the element-wise means and standard deviations. Acoustic signals in the CLEAR dataset have variable duration. To work around this constraint, we padded the spectrograms with zeros so that they all have the same dimensions. Another solution is to resize all spectrograms to a fixed size using bilinear interpolation as it is commonly done in the visual domain. From a purely acoustic point of view, we hypothesized that padding spectrograms would preserve the relative time axis between all scenes and would yield better results than resizing. We revisited this hypothesis by training our best architecture on resized spectrograms. Surprisingly, the network performed better by about 1 percentage points when trained on the resized spectrograms. We do not have a definitive answer as to why we observe this behavior. It might be due to the fact that padded spectrograms contain less information since a portion of them is filled with zeros. The network could also be using the different time resolutions in the resized spectrograms as a cue to cheat the reasoning process. However, this could also happen when the scenes are padded since the padded section is unique to each scene. We analyzed the network activations and did not find that the padded section was overly activated. We leave it up to future work to further analyze this.
Supplementary F Importance of text versus audio modality
To validate the fact that NAAQA makes good use of both input modalities (spectrogram and text), the network was trained using only one modality. To do so, either all values of the input spectrograms were set to 1 or all questions were set to the unknown token during training. No more than 48% of accuracy on the test set was achieved for either of these scenarios which indicates that the network uses information from both modalities to achieve high performance and that the dataset cannot be solved using only one modality with this architecture. This is however much higher than the probability of answering the correct answer by chance ($\frac{1}{57}=1.75\%$), or by majority vote (7.5%) introduced in Section 3.1. Although we have carefully designed the data set in order to avoid bias, some form of bias is intrinsic in QA problems. For example, an acoustic sequence containing only $n$ elementary sounds automatically disables all counting answers above $n$ regardless which question is asked. Similarly, asking a yes/no question limits the number of possible answers to two, regardless of the acoustic scene.
Supplementary G DAQA’ missing sounds
Here is a list of all the missing elementary sounds that were originally in DAQA and that are not in the DAQA’ dataset (because of not being anymore accessible on youtube) :
b000_20, b000_3, b000_4, b000_5, b000_6, b001_11, c001_13, c001_14, c001_15, c001_16, c001_17, c001_18, c001_19, c001_20, c003_9, c004_10, d001_1, d001_9, d002_19, f001_19, h001_10, h001_11, h001_12, h001_13, h001_14, h001_15, h001_19, h001_2, h001_3, h001_4, h001_5, h001_6, h001_7, h001_8, h001_9, h002_10, h002_11, h002_12, h002_8, h002_9, t000_14, t000_15
Supplementary H Network Optimization
Gru (G) Overall Accuracy (%) Nb of
Units Train Val Test Parameters
512 95.5 ±0.28 90.1 ±0.37 79.6 ±0.30 1.94 M
256 94.3 ±0.51 90.5 ±0.30 79.5 ±0.05 1.68 M
1,024 94.2 ±0.31 89.7 ±0.18 79.0 ±0.14 2.46 M
4,096 94.8 ±0.25 89.9 ±0.25 78.9 ±0.23 5.61 M
2,048 95.8 ±0.14 90.0 ±0.05 78.7 ±0.04 3.51 M
TABLE VIII: Impact of the number of GRU text-processing units $G$. The interleaved extractor is used with hyper-parameters from the initial configuration described in Section 5.3 except for $G$. The number of GRU unit has a strong impact on the number of parameters. Setting $G=512$ actually increases accuracy while reducing the number of parameters by a factor of $\sim$2.8. The rows are ordered by test accuracy.
Nb (J) Nb Filter (M) Overall Accuracy (%) Nb of
Resblock in Resblock Train val Test Parameters
4 128 94.9 ±0.91 89.9 ±0.39 79.1 ±0.56 1.94 M
3 128 95.0 ±0.67 89.4 ±0.42 78.4 ±0.25 1.64 M
2 128 93.6 ±0.40 86.9 ±0.61 77.3 ±0.88 1.35 M
4 64 92.8 ±0.95 87.6 ±0.98 77.0 ±0.66 1.11 M
3 64 91.8 ±1.36 86.2 ±0.77 76.4 ±0.65 1.01 M
2 64 90.8 ±0.75 84.5 ±0.64 75.0 ±0.63 0.90 M
4 32 88.1 ±1.35 83.1 ±1.08 73.6 ±1.18 0.82 M
3 32 86.3 ±1.39 80.6 ±1.54 71.5 ±1.54 0.78 M
2 32 84.4 ±1.36 76.8 ±2.02 68.4 ±1.59 0.74 M
1 128 89.0 ±1.80 75.3 ±2.65 67.2 ±2.43 1.05 M
1 64 84.3 ±4.61 69.6 ±4.61 62.1 ±3.41 0.79 M
1 32 67.0 ±3.58 57.3 ±2.07 52.9 ±1.50 0.69 M
TABLE IX: Impact of reducing the number of Resblock $J$ and the number of filters $M$ in each Resblocks. The interleaved extractor is used with hyper-parameters from the initial configuration described in Section 5.3 except for $J$, $M$, $C=256$, $H=1024$ and $G=512$. With $J=4$ and $M=64$, the network reaches great accuracy with only 0.67M parameters (vs 5.61M for the initial configuration). The accuracy drops when $J<3$ and $M=32$. The configuration $J=3$ and $M=128$ get stuck in local minimums often which explain the large STD. The rows are ordered by test accuracy
Tables VIII and IX report the intermediate results of the NAAQA network complexity study on CLEAR2.